# GHOST: Hallucination-Inducing Image Generation for Multimodal LLMs

### v1

@@ 6.2 Mitigation
As a small toy setup, we fine-tune Qwen2.5-VL using GHOST images. Following DASH (Augustin et al., 2025), we include both positive and negative samples to avoid forgetting of object concepts. However, unlike DASH, which uses real COCO images as positives, we generate both positives and negatives synthetically using Diffusion unCLIP to ensure consistency in visual style. This prevents the model from simply learning to reject all synthetic content.
Table 4 shows that fine-tuning improves robustness to hallucination-inducing images generated by other models (LLaVA and GLM), despite never seeing them during training. We also observe improvements on the POPE (Li et al., 2023) benchmark, while DASH reports no gains on POPE. Finally, performance on general tasks such as VQAv2 and COCO captioning remains nearly unchanged, suggesting that fine-tuning likely preserves general capabilities.
@@
GHOST-LLaVA GHOST-GLM POPE VQAv2 Caption
Baseline 52.6 63.1 88.5 73.5 89.5
Finetuned 7.0 7.3 90.7 73.4 89.4
@@ 7 Conclusion
We introduce GHOST, a pipeline for generating images that induce hallucinations in MLLMs by inserting minimal cues. Our method leverages diffusion models to craft these cues, demonstrating that their vision space aligns well with the visual representations used by MLLMs. Beyond a single model, GHOST reveals cross-modal failure modes that are not detectable by prior works. We hope GHOST can help and inspire future work toward building more robust and reliable multimodal systems.
Acknowledgments
This project was supported in part by a grant from an NSF CAREER AWARD 1942230, the ONR PECASE grant N00014-25-1-2378, ARO’s Early Career Program Award 310902-00001, Army Grant No. W911NF2120076, the NSF award CCF2212458, NSF Award No. 2229885 (NSF Institute for Trustworthy AI in Law and Society, TRAILS), a MURI grant 14262683, DARPA AIQ DARPA AIQ grant HR00112590066 and an award from meta 314593-00001. Naveed Akhtar is a recipient of the Australian Research Council Discovery Early Career Researcher Award (project # DE230101058) funded by the Australian Government. This work is also partially supported by Google Research Scholar Program Award.
Appendix A Human Evaluation
@@
Appendix B DASH
We reproduce DASH results with Qwen2.5 and LLaVA. For LLaVA, we only evaluate DASH-LLM, a variant of DASH that relies solely on language model feedback. We omit DASH-OPT as it is too memory-intensive, even a batch size of 1 exceeds the limits of 80GB GPUs. For Qwen2.5, we evaluate both DASH-LLM and DASH-OPT. In DASH-OPT, we tune three parameters: step size, gradient clipping, and the VLM loss weight. In DASH-OPT, we tune three parameters: step size, gradient clipping, and the VLM loss weight. Specifically, we search step size and gradient clip values in $[0.1,0.2,0.4,0.8]$, and the VLM loss weight in $[1,2,4,6,8]$. As our evaluation metric, we count the number of images that do not contain the target object according to an object detector but still mislead the MLLM into hallucinating its presence. We are unable to use ReLION-5B, as its KNN search service is private. Additionally, computing CLIP embeddings over the full dataset is computationally infeasible for us. Instead, we use the COCO dataset for our experiments. The corresponding results are shown below.
@@ Appendix B DASH
Appendix C Mapper
C.1 Mapper Training
@@
C.3 Mapper Evaluation
Inspired by (Liu et al., 2023), we assess the reconstruction capability of the MLP connector $f$ using GPT-4 as a judge. For each input image $x$, we first obtain the response of the target MLLM (e.g., LLaVA) when conditioned on the true image embedding $\mathcal{V}_{\mathcal{M}})$. We then provide this response, together with the ground-truth COCO object annotations, to GPT-4, which rates the consistency between the response and the annotations. Next, we replace the image input with the projected embedding $\Pi(\mathcal{V}_{\text{clip}}(x))$, obtain the MLLM response, and request GPT-4 to rate it in the same way. The relative score between the two ratings serves as a measure of the mapper’s ability to reconstruct the semantics of the image. We define the relative score as the ratio between the GPT rating of the projected-embedding response and that of the real-image response which quantifies how much semantic information from the original image embedding is retained after projection. Further details on this experiment and the evaluation prompt are provided in the appendix. For this study, we randomly selected 100 images from the COCO dataset. The results on LLaVA-v1.6, Qwen2.5-VL, and GLM are reported in Table 11. As shown, the projector preserves most of the semantic content of the original image while substantially reducing computational cost.
Table 11: Evaluation of Mapper using GPT-4 as judge. Scores are reported for MLLM responses conditioned on the real image and on the reconstructed embedding.
@@ C.3 Mapper Evaluation
Appendix D Latent Diffusion Models
@@
Appendix E GHOST Optimization
E.1 GHOST Without CLIP Sorting
For evaluating the performance of GHOST without CLIP-based sorting, we conduct an experiment on four classes: “Boat”, “Vase”, “Knife”, and “Bottle”. For each class, we randomly select 200 samples that do not have the target object from their categories based on COCO annotations. When we select the pictures without sorting, the initial images are farther from the target object semantically, so more changes are needed to induce hallucination. To overcome this issue, we use noise level $t=15$ (in contrast to our standard setting where the noise level is $t=30$) as a hyperparameter to add more noise to the image and be able to change the image further. We also select $\tau=0.6$ to soften the constraint of optimization; all the other hyperparameters are the same as the standard setting. As reported in Table 12, the success rate is comparable to our method in the sorting setting. Qualitative samples with $t=15$ as the noise level are also shown in Figure 12. As shown in Figure 12, when the noise level is decreased, more changes are applied to the initial images to induce hallucination.
@@ E.1 GHOST Without CLIP Sorting
Table 12: Class-wise results with success rate and number of successful samples in $t=15$ noise level setting.
Class Images Optimized Success Rate Successful Samples
Vase 198 0.32 63
Knife 198 0.46 91
Boat 200 0.24 48
Bottle 188 0.57 107
Total / Mean 784 0.40 309
E.2 Text representation
@@ E.4 Prompts
E.5 Effect of $\lambda_{reg}$
To assess the effect of $\lambda_{reg}$, we apply GHOST to the GLM4.1 Thinking model with $\lambda_{reg}\in\{1.0,1.5,2.0\}$. Since the FID score is sensitive to the number of samples, for each pair of $\lambda_{reg}$ values we select only the samples that are successful in both settings, and then compute the FID score against both the COCO validation set and the original images. As reported in Table 15, increasing $\lambda_{reg}$ consistently makes the generated images more similar to the original ones, leading to improved realism, as reflected by lower FID scores.
E.6 Effect of $\lambda_{clip}$
Increasing $\lambda_{\text{clip}}$ encourages semantic separation between the guiding embedding $z$ and the target-object semantics. We implement GHOST on two target objects (“traffic light" and “Boat", 200 images each) on Qwen2.5-VL with $\lambda_{\text{clip}}\in\{5,10,15,20\}$ and N=4N{=}4 samples per image to evaluate the effect of it. As shown in Fig. 14, larger $\lambda_{\text{clip}}$ consistently reduces the count of OD-filtered samples, indicating that the regularization term achieves its intended effect by suppressing target-object semantics in the generated images. From $\lambda_{\text{clip}}=5$ to $15$, stronger suppression of target-object cues makes it harder for the MLLM to answer “yes”, so more samples are generated and also a larger share of generated samples fail to trigger hallucination. At $\lambda_{\text{clip}}=20$, this term becomes over-weighted: fewer runs satisfy $p(y^{*}\mid\mathbf{X}_{q},\Pi(z))\geq\tau_{\text{yes}}$, and more runs hit the maximum (100) steps of optimization (e.g., $10\rightarrow 20$)
E.7 Effect of $\tau$
We implement GHOST with two target objects (200 images each) on LLaVA-1.6, varying $\tau\in\{0.5,0.6,0.7,0.8,0.9\}$ and using N=4N{=}4 generations per image. As shown in Fig. 14, increasing $\tau$ tightens the optimization constraints, and the number of generated samples consistently decreases. At lower $\tau$, the embedding contains weaker target-object cues, so more generations fail to elicit a “yes” from the MLLM. For $\tau\in\{0.5,0.6,0.7\}$, the ratio of total to unsuccessful generations is similar, and the number of successful samples is nearly unchanged.
@@ E.7 Effect of $\tau$
Appendix F GHOST Additional Results

### v2

@@ 6.2 Mitigation
As a small toy setup, we fine-tune Qwen2.5-VL with LoRA (Hu et al., 2022) using GHOST images. Following DASH (Augustin et al., 2025), we include both positive and negative samples to avoid forgetting of object concepts. However, unlike DASH, which uses real COCO images as positives, we generate both positives and negatives synthetically using Diffusion unCLIP to ensure consistency in visual style. This prevents the model from simply learning to reject all synthetic content.
Table 4 shows that fine-tuning Qwen2.5-VL on GHOST images improves robustness to hallucination. We observe consistent gains on POPE (Li et al., 2023) as well as on the CHAIR benchmark, following the evaluation setup from Li et al. (2023). Also, performance on general vision-language tasks (VQAv2 and COCO Captioning) remains nearly unchanged, suggesting that GHOST-based fine-tuning preserves general capabilities. Additional fine-tuning details can be found in Appendix J.
@@
POPE $\uparrow$ CHAIRs $\downarrow$ CHAIRi $\downarrow$ VQAv2 $\uparrow$ Caption $\uparrow$
Baseline 88.7 3.8 3 89.5 72.8
Finetuned 93.2 2.9 2.6 89.4 71.5
Figure 10: Applying GHOST to other types of hallucination.
6.3 Extending GHOST
GHOST primarily targets object hallucinations. However, to demonstrate its broader potential, we conduct a small experiment to test whether GHOST can also stress-test other types of hallucinations, such as attribute or relation hallucinations. Using the same losses as in the main setting, we select samples for which the victim model originally answers “No” to five queries, including “someone is holding a knife,” “the vase is transparent,” “the bottle is made of glass,” “the clock is hanging on the wall,” and the relation query “the boat is near the shore.” We then optimize the embedding to flip the first-token prediction to “Yes.” No object detector is used in this setting, as these queries do not correspond to object presence. Qualitative samples are shown in Figure 10. Across these five attribute/relation queries, GHOST achieves an average success rate of 32.2%. More details are provided in Appendix H.
7 Limitation & Future work
A key practical limitation of GHOST is its reliance on diffusion models that support image embedding-level conditioning, such as unCLIP. GHOST optimizes directly in the vision-embedding space, which requires a generative model capable of decoding modified embeddings back into pixel space. Many popular diffusion architectures do not currently provide this interface and therefore cannot serve as drop-in replacements in our framework. In addition, our current evaluation primarily focuses on object-centric hallucinations; a more systematic investigation across a broader range of hallucination modalities, along with more extensive corrective evaluations across architectures and hallucination types, would provide a more comprehensive assessment.
@@ 8 Conclusion
We introduce GHOST, a pipeline for generating images that induce hallucinations in MLLMs by inserting misleading cues. Our method leverages diffusion models to craft these cues, demonstrating that their vision space aligns well with the visual representations used by MLLMs. Beyond a single model, GHOST reveals cross-modal failure modes. We hope GHOST can help and inspire future work toward building more robust and reliable multimodal systems.
Reproducibility. We provide detailed descriptions of all components to facilitate reproducibility. Appendix D outlines the training procedure for the mapper, Appendix F describes the GHOST pipeline, and Appendix J covers the fine-tuning setup and evaluation. We have also submitted our code as supplementary material.
Appendix A AI Usage Clarification
Large Language Models were used to to aid with writing clarity and polish.
Appendix B Human Evaluation
@@
Naturalness. The goal of GHOST is not to produce perfectly photorealistic images, but to create semantically meaningful perturbations that reveal model vulnerabilities. While FID scores in the main paper show that GHOST preserves the quality of the underlying diffusion model, we additionally assess naturalness with a human study. We sample 30 GHOST images and 30 unCLIP-generated counterparts guided by the same initial images. Volunteers rate naturalness on a 1–5 scale (1: completely unnatural, 5: completely natural). As shown in Table 6, GHOST images receive ratings comparable to the diffusion baseline (3.55 vs. 3.56 on average), with a similar proportion rated “mostly natural” (ratings $\geq 4$).
Table 6: Human Evaluation for Naturalness of GHOST Images.
Model Avg. Rating # Votes # $\geq$ 4
GHOST 3.55 191 111
SD-unCLIP 3.56 189 118
Appendix C DASH
We reproduced DASH results with Qwen2.5 and Llava. For Llava we only investigated DASH-LLM which is a varient of DASH results. We omited DASH-OPT since even batch size of 1 could not be achieved with 80GB gpus. For Qwen2.5 we investigated both DASH-LLM and DASH-OPT. For DASH-OPT we finetuned three parameters, step size, gradient clip and vlm weight on the loss. for step size and gradient clip we investigated $[0.1,0.2,0.4,0.8]$ and for MLLM weight we checked $[1,2,4,6,8]$. Our meassure was the number images that does not contain the object according to object detector but mislead the MLLM. We could not use ReLION-5B as the KNN search service is private. In addition calculating clip embeddings on all those images was not practical for us. We used COCO dataset instead. On COCO these are the results.
@@ Appendix C DASH
Appendix D Mapper
D.1 Mapper Training
@@
D.3 Mapper Evaluation
Inspired by (Liu et al., 2023a), we assess the reconstruction capability of the MLP connector $f$ using GPT-4 as a judge. For each input image $x$, we first obtain the response of the target MLLM (e.g., LLaVA) when conditioned on the true image embedding $\mathcal{M}(\mathcal{E}_{v}(x))$. We then provide this response, together with the ground-truth COCO object annotations, to GPT-4, which rates the consistency between the response and the annotations. Next, we replace the image input with the projected embedding $f(\mathcal{E}_{\text{clip}}(x))$, obtain the MLLM response, and request GPT-4 to rate it in the same way. The relative score between the two ratings serves as a measure of the mapper’s ability to reconstruct the semantics of the image. We define the relative score as the ratio between the GPT rating of the projected-embedding response and that of the real-image response which quantifies how much semantic information from the original image embedding is retained after projection. Further details on this experiment and the evaluation prompt are provided in the appendix. For this study, we randomly selected 100 images from the COCO dataset. The results on LLaVA-v1.6, Qwen2.5-VL, and GLM are reported in Table 12. As shown, the projector preserves most of the semantic content of the original image while substantially reducing computational cost.
Table 12: Evaluation of projector reconstruction using GPT-4 as judge. Scores are reported for MLLM responses conditioned on the real image and on the reconstructed embedding.
@@ D.3 Mapper Evaluation
Appendix E Latent Diffusion Models
@@
Appendix F GHOST Optimization
F.1 GHOST Without CLIP Sorting
For evaluating the performance of GHOST without CLIP-based sorting, we conducted an experiment on four classes: “Boat”, “Vase”, “Knife”, and “Bottle”. For each class, we randomly selected 200 samples that do not have the target object from their categories based on COCO annotations. When we select the pictures without sorting, the initial images are farther from the target object semantically, so more changes are needed to induce hallucination. To overcome this issue, we used noise level $t=15$ (in contrast to our standard setting where the noise level is $t=30$) as a hyperparameter to add more noise to the image and be able to change the image further. We also selected $\tau=0.6$ to soften the constraint of optimization; all the other hyperparameters are the same as the standard setting. As reported in Table 13, the success rate is comparable to our method in the sorting setting. Qualitative samples with $t=15$ as the noise level are also shown in Figure 13. As shown in Figure 13, when the noise level is decreased, more changes are applied to the initial images to induce hallucination.
@@ F.1 GHOST Without CLIP Sorting
Table 13: Class-wise hallucination results at noise level $t=15$.
Class Samples Hallucination Success (%)
Vase 198 63 32%
Knife 198 91 46%
Boat 200 48 24%
Bottle 188 107 57%
Total / Mean 784 309 40%
F.2 Text representation
@@ F.4 Prompts
F.5 Effect of $\lambda_{reg}$
To assess the effect of $\lambda_{reg}$, we applied GHOST to the GLM4.1 Thinking model with $\lambda_{reg}\in\{1.0,1.5,2.0\}$. Since the FID score is sensitive to the number of samples, for each pair of $\lambda_{reg}$ values we selected only the samples that were successful in both settings, and then computed the FID score against both the COCO validation set and the original images. As reported in Table 16, increasing $\lambda_{reg}$ consistently makes the generated images more similar to the original ones, leading to improved realism, as reflected by lower FID scores.
F.6 Effect of $\lambda_{clip}$
Increasing $\lambda_{\text{clip}}$ encourages semantic separation between the guiding embedding $z$ and the target-object semantics. We implemented GHOST on two target objects (“traffic light” and “Boat”, 200 images each) on Qwen2.5-VL with $\lambda_{\text{clip}}\in\{5,10,15,20\}$ and N=4N{=}4 samples per image to evaluate the effect of it. As shown in Fig. 15, larger $\lambda_{\text{clip}}$ consistently reduces the count of OD-filtered samples, indicating that the regularization term achieves its intended effect by suppressing target-object semantics in the generated images. From $\lambda_{\text{clip}}=5$ to $15$, stronger suppression of target-object cues makes it harder for the MLLM to answer “yes”, so more samples are generated and also a larger share of generated samples fail to trigger hallucination. At $\lambda_{\text{clip}}=20$, this term becomes over-weighted: fewer runs satisfy $p(y^{*}\mid\mathbf{X}_{q},\Pi(z))\geq\tau_{\text{yes}}$, and more runs hit the maximum (100) steps of optimization (e.g., $10\rightarrow 20$)
F.7 Effect of $\tau$
We implemented GHOST with two target objects (200 images each) on LLaVA-1.6, varying $\tau\in\{0.5,0.6,0.7,0.8,0.9\}$ and using N=4N{=}4 generations per image. As shown in Fig. 15, increasing $\tau$ tightens the optimization constraints, and the number of generated samples consistently decreases. At lower $\tau$, the embedding contains weaker target-object cues, so more generations fail to elicit a “yes” from the MLLM. For $\tau\in\{0.5,0.6,0.7\}$, the ratio of total to unsuccessful generations is similar, and the number of successful samples is nearly unchanged.
@@ F.7 Effect of $\tau$
F.8 Runtime Analysis
[…]
Table 17: Runtime of GHOST across six objects on a single A100 GPU for Qwen2.5-VL.
Category Samples Avg (s) Min (s) Max (s) Total (min)
Boat 378 8.59 2.23 20.04 56.20
Bottle 364 9.73 2.24 19.92 61.08
Carrot 387 9.97 2.13 20.04 66.37
Vase 380 10.40 2.22 20.19 67.90
Suitcase 380 9.02 2.27 20.74 59.21
Traffic Light 398 7.18 2.24 19.60 49.66
F.9 Structural Similarity Index (SSIM)
To further quantify perceptual similarity and assess the preservation of semantic content, we compute the average and standard deviation of SSIM (Wang et al., 2004) across GHOST, unCLIP, and Stable Diffusion outputs using the same setup as our FID semantic fidelity experiment. Higher SSIM indicates greater structural similarity. We obtain:
  • GHOST: $0.6051\pm 0.1156$
  • unCLIP: $0.2322\pm 0.1218$
  • Stable Diffusion: $0.1843\pm 0.1041$
These results indicate that GHOST samples preserve the structural and semantic content of the original images more strongly than standard diffusion outputs.
F.10 Higher Resolution Scalability
To evaluate whether GHOST scales to higher-resolution inputs, we increase the image resolution from $\times336$ (the baseline used in the main paper) to $\times756$. We train a separate mapper for the higher resolution and run GHOST on two object categories, traffic light and boat, using 200 samples each on an A100 GPU with Qwen2.5-VL-7B as the victim model. Table 18 reports the runtime statistics and attack success rates.
Across both categories, GHOST continues to run efficiently at the higher resolution. The per-image runtime increases moderately, as expected due to the larger number of visual tokens, while the success rates remain comparable to the $\times336$ baseline. These results indicate that GHOST scales to higher resolutions without becoming prohibitively slow. Further speed-ups are possible by parallelizing optimization across samples or adjusting hyperparameters such as the number of optimization steps and $N$.
Table 18: GHOST results on $\times756$ resolution input.
Category Samples Avg / Sample (s) Min (s) Max (s) Total (min) ASR
Traffic Light 249 16.50 2.54 37.12 70.64 0.33
Boat 295 9.77 2.49 37.80 50.15 0.32
Appendix G GHOST on ObjectNet
To evaluate GHOST beyond COCO, we apply our method to 10 categories from ObjectNet (Barbu et al., 2019), using Qwen2.5-VL-7B as the victim model. For each category, we optimize 400 samples, excluding cases where the model already hallucinates the target object (consistent with the main experiments). Table 19 reports the success rates, number of optimized samples, and number of successful samples for all categories. GHOST achieves an overall success of 28%, which is comparable to our results on COCO. Qualitative examples from ObjectNet are shown in Figure 16.
Figure 16: GHOST examples on ObjectNet dataset.
Table 19: GHOST results on 10 ObjectNet categories.
Category Samples Hallucination Success (%) Images Generated
Bracelet 312 100 32% 555
Cellphone 336 134 40% 1017
Helmet 377 124 33% 1259
Lemon 346 80 23% 1205
Microwave 313 38 12% 1128
Pillow 216 102 47% 684
Printer 357 82 23% 1244
Watch 397 67 17% 1244
Wallet 293 141 48% 877
Webcam 385 50 13% 1294
Appendix H Extending GHOST
GHOST primarily targets object hallucinations. However, to demonstrate its broader potential, we conduct a small experiment to test whether GHOST can also stress-test other types of hallucination, such as attribute or relation hallucinations. Using the same losses as in the main setting, we select samples for which the victim model originally answers “No” to five queries, including “someone is holding a knife,” “the vase is transparent,” “the bottle is made of glass,” “the clock is hanging on the wall,” and the relation query “the boat is near the shore.” We then optimize the embedding to flip the first-token prediction to “Yes.” No object detector is used in this setting, as these queries do not correspond to object presence.
Table 20 summarizes the five queries evaluated, along with the number of selected samples and the final success rates. The average success rate of 32.2% is comparable to the object-existence success rate (29.9%) reported in the main paper. These results provide preliminary evidence that GHOST can extend beyond object-centric hallucinations to induce attribute-level and relation-level failures.
Table 20: Attribute and relation hallucination experiments. Each query contains only samples for which the model originally answered “No”.
Query Type Example Question # Initial Samples Success
Relation Is someone holding a knife? 100 16%
Relation Is the clock hanging on the wall? 74 36%
Attribute Is the vase transparent? 93 39%
Attribute Is the bottle made of glass? 67 15%
Relation Is the boat near the shore? 53 55%
Average 32.2%
Hyperparameter Settings. Table 21 lists the exact optimization hyperparameters used for each query.
Table 21: GHOST hyperparameters for attribute and relation hallucination experiments.
Query lr total steps $\tau$ N Guidance Scale $\lambda_{clip}$ $\lambda_{reg}$ OD thr. t num_inf.
Boat near shore 0.1 100 0.8 4 7.5 5.0 5.0 0.5 30 50
Someone holding a knife 0.1 100 0.8 4 7.5 5.0 5.0 0.5 30 50
Bottle made of glass 0.1 100 0.8 4 7.5 15.0 10.0 0.5 40 50
Clock hanging on wall 0.1 100 0.8 4 7.5 5.0 5.0 0.5 30 50
Vase is transparent 0.1 100 0.8 4 7.5 15.0 10.0 0.5 30 50
# Attribute templates
attribute_templates = [
"Is the {obj} {attr}? Answer with Yes or No’.",
"Is the {obj} {attr} in the image? Answer with Yes or No’.",
"Would you say the {obj} is {attr}? Answer with Yes or No’.",
"Can you confirm the {obj} is {attr}? Answer with Yes or No’.",
"Does the {obj} appear {attr}? Answer with Yes or No’.",]
# Relation templates
relation_templates = [
"{rel}? Answer with Yes or No’.",
"Is it true that {rel}? Answer with Yes or No’.",
"Can you confirm that {rel}? Answer with Yes or No’.",
"Does the image show that {rel}? Answer with Yes or No’.",
"Would you say that {rel}? Answer with Yes or No’.",]
Listing 3: Prompt templates used for attribute-based and relation-based hallucination experiments. The placeholders ‘obj’, ‘attr’, and ‘rel’ are filled with the corresponding object name, attribute phrase, or relational statement during optimization.
Prompt Templates. We use two sets of prompt templates: one for attributes and one for relations. These templates are shown in Listing 3.
Appendix I Object Detector Usage
In GHOST, the object detector serves only as a sanity-check filter to remove images that may still contain the target object. The detector does not participate in the optimization process, all losses operate entirely in the CLIP embedding space and are designed to remove object-specific information. To ensure that detector false negatives do not impact our results, we additionally perform a human evaluation confirming that the retained samples do not contain the target object (Sec. 5.1). Since the optimization itself is detector-agnostic, GHOST does not depend on any particular detector family.
Ensemble experiment. To further assess detector dependence, we apply a strict ensemble filter using four detectors spanning both open-vocabulary and closed-set families: OWLv2 (Minderer et al., 2023), GroundingDINO (Liu et al., 2023b), YOLO-World (Cheng et al., 2024), and YOLOv8 (Jocher et al., 2023). A generated sample is discarded if any detector assigns confidence $\geq 0.5$ to the target object. We evaluate 300 samples across five object classes using Qwen2.5-VL as the victim model. The results are reported in Table 22: even under this conservative filtering, GHOST maintains a 22.7% success across 1,419 images.
Table 22: Ensemble of Object Detectors: A sample is removed if any of four detectors predict the target object with confidence $\geq 0.5$.
Object Success Samples Hallucination
Boat 16% 279 45
Bottle 25% 271 68
Clock 15% 290 44
Suitcase 29% 280 81
Traffic Light 28% 299 84
Overall 22.7% 1419 322
Appendix J Finetuning
Setup. We fine-tune Qwen2.5-VL-7B on its GHOST-generated images to assess whether such counterfactual samples improve model performance. This is intended as a proof-of-concept demonstration, with large-scale training and full-class coverage left for future work. We use LoRA (Hu et al., 2022) for parameter-efficient adaptation, BitsAndBytes (Dettmers et al., 2023) for quantization, and TRL (von Werra et al., 2020) for SFT implementation. All experiments are conducted on an 8-way L40S GPU node. Full configuration details are provided in Table 23.
Dataset. We construct a balanced fine-tuning set by sampling 150 negative images (GHOST images) and generating 150 positive images for each class, resulting in 3,000 samples in total. Positive samples are synthesized using Stable Diffusion unCLIP (Rombach et al., 2022), conditioned on images that contain the target object. We avoid using real COCO images as positives so that the model cannot rely on a trivial shortcut such as “synthetic = negative, real = positive.” We use the prompt: Is there a {obj} in the image? Answer with ’Yes’ or ’No’. with “Yes” for positives and “No” for negatives.
To increase label fidelity for the synthesized positives, we rely on COCO annotations and filter candidates by ranking their CLIP embeddings against a compositional text embedding of the target object (Sec. F.2), keeping the highest-scoring images as inputs to the unCLIP model. This improves the likelihood that the positive set truly depicts the intended object while keeping the data domain consistent (synthetic vs. synthetic). Example positive samples are shown in Fig. 17.
Model Selection. Since LoRA fine-tuning can be sensitive to hyperparameters, we train multiple configurations and select the best-performing model on POPE (that is reported in Table 23). We generally observe that the rank, learning rate, and batch size are the most influential factors for effective fine-tuning.
[…]
To assess the impact of fine-tuning on overall model ability, we evaluate the model on both VQA and image captioning. For VQA, we use the small VQAv2 subset from Hugging Face [^1] https://huggingface.co/datasets/merve/vqav2-small, which contains image–question pairs with a single-word ground-truth answer. We report accuracy for both the base and fine-tuned models. For captioning, we use the COCO validation images and their annotations as ground truth. We generate captions using the prompt “Write a short caption for the given image.” for both models and compute BERTScore (Zhang* et al., 2020) between the generated captions and COCO references. These evaluations measure whether GHOST fine-tuning impacts general capabilities beyond hallucination robustness.
Table 23: Fine-tuning configuration for Qwen2.5-VL.
Quantization (BitsAndBytes)
4-bit Quantization NF4
Compute Dtype bfloat16
Quant Storage bfloat16
Double Quant True
LoRA Configuration
Rank ($r$) 2
Alpha 2
Dropout 0.05
Target Modules all-linear
Modules Saved lm_head, embed_tokens
Training Setup (SFTConfig)
Optimizer AdamW (fused)
Learning Rate $1\times 10^{-5}$
Epochs 6
Batch Size 8
Grad. Accumulation 1
Precision bf16
Seed 42
Appendix K GHOST Additional Results